# The variational principle for a marked Gibbs point process with infinite-range multibody interactions

### v1

@@ 8.3. Tempered boundary conditions sampled from a finite-energy measure
As a last step, we now show how to extend the existence of the pressure to boundary conditions sampled from a tempered measure $P$ with finite energy.
@@ 9.2. Gibbs measures are minimizers
We now show that infinite-volume Gibbs measures are minimizers of the variational formula. For this, we first show that the energy density does not depend on the choice of boundary conditions. This will allow us to restrictions of infinite-volume Gibbs measures to the finite-volume Gibbs measure with free boundary conditions.
@@
$=\beta P\left[H_{\Lambda,\emptyset}(\bullet)-H_{\Lambda}(\bullet)\right]+\left[\log(Z_{\Lambda,z,\beta})-P(\log(Z_{\Lambda,z,\beta,\bullet}))\right].$
Moreover, by Fatou’s lemma
@@
[…]
@@
$\limsup_{n\to\infty}\frac{I(P_{\Lambda_{n}}\,|\,G_{\Lambda_{n},z,\beta})}{|\Lambda_{n}|}\leq 0\leq\liminf_{n\to\infty}\frac{I(P_{\Lambda_{n}}\,|\,G_{\Lambda_{n},z,\beta})}{|\Lambda_{n}|}$
which of course implies
@@ 9.3. Minimizers are Gibbs measures
$\left\|G_{L,z,\beta,\zeta_{\Lambda_{k}}}-G_{L,z,\beta,\zeta}\right\|_{\text{TV}}\leq 2\frac{\pi^{z}_{L}\left(|\mathrm{e}^{-\beta H_{L,\zeta}}-\mathrm{e}^{-\beta H_{L,\zeta_{\Lambda_{k}}}}|\right)}{\max\big{\{}Z_{L,z,\beta,\zeta},Z_{L,z,\beta,\zeta_{\Lambda_{k}}}\big{\}}}\leq 4\mathrm{e}^{z\left\lvert L\right\rvert}\,\pi^{z}_{L}\left(B_{K}^{c}\right),$
@@
Now note that, if $0=\delta F(P)=\lim_{n\to\infty}\frac{1}{\left\lvert\Lambda_{n}\right\rvert}I(P_{\Lambda_{n}}\,|\,G_{\Lambda_{n},z,\beta})$, then, by the translation invariance of $P$, it follows that $P_{\Lambda}\ll G_{\Lambda,z,\beta}$ with some density $g_{\Lambda}$ for large enough cubes $\Lambda$. For every subset $\Delta\subset\Lambda$, we consider the restriction
@@
Furthermore,
@@
Using translation invariance (i.e.,$I_{\Delta+x}(P\,|\,G_{\Lambda+x,z,\beta})=I_{\Delta}(P\,|\,G_{\Lambda,z,\beta})$), we have thus found suitable sets $\Lambda:=\Lambda_{n}-i(l)$ and $\Delta:=\Delta(l)-i(l)$. ∎
@@
$\int\,P(\mathrm{d}\zeta)\,f(\zeta)$ =$\stackrel{{\scriptstyle(1)}}{{=}}\int\,G_{\Lambda,z,\beta}(\mathrm{d}\zeta)\,g_{\Lambda,\Delta}(\zeta)\,f(\zeta)$
@@
=$\stackrel{{\scriptstyle(3)}}{{=}}\int\,G_{\Lambda,z,\beta}(\mathrm{d}\zeta)\,g_{\Lambda,\Delta\setminus L}(\zeta)\,\left[\int\,G_{L,z,\beta,\zeta}(\mathrm{d}\omega)\,f(\omega_{L}\zeta_{L^{c}})\right]$
@@
=$\stackrel{{\scriptstyle(5)}}{{=}}\int\,P(\mathrm{d}\zeta)\,\left[\int\,G_{L,z,\beta,\zeta_{\Lambda_{k}}}(\mathrm{d}\omega)\,f(\omega_{L}\zeta_{\Lambda_{k}\setminus L})\right]$

### v2

@@ 8.3. Tempered boundary conditions sampled from a finite-energy measure
As a last step, we now show how to extend the existence of the pressure to boundary conditions sampled from a tempered measure $P$ with finite energy density. These configurations will then of course always respect the hardcore condition.
@@ 9.2. Gibbs measures are minimizers
We now show that infinite-volume translation-invariant Gibbs measures are minimizers of the variational formula. For this, we first show that the energy density does not depend on the choice of boundary conditions. This will allow us to compare restrictions of infinite-volume Gibbs measures to the finite-volume Gibbs measure with free boundary conditions.
@@
$=\beta P\left[H_{\Lambda,\emptyset}(\bullet)-H_{\Lambda}(\bullet)\right]+\left[\log(Z_{\Lambda,z,\beta})-P(\log(Z_{\Lambda,z,\beta,\bullet}))\right],$
where we used the notation $Q_{2}^{\omega}\left[Q_{1}^{\zeta}\left[f(\omega,\zeta)\right]\right]=\int\,Q_{2}(\mathrm{d}\omega)\int\,Q_{1}(\mathrm{d}\zeta)f(\omega,\zeta)$ for readability reasons. Moreover, by Fatou’s lemma
@@
$\leq\beta\limsup_{n\to\infty}\frac{P\left[H_{\Lambda_{n},\emptyset}(\bullet)-H_{\Lambda_{n}}(\bullet)\right]}{|\Lambda_{n}|}+\left[\limsup_{n\to\infty}\frac{\log(Z_{\Lambda_{n},z,\beta})}{|\Lambda_{n}|}-\liminf_{n\to\infty}\frac{P[\log(Z_{\Lambda_{n},z,\beta,\bullet})]}{|\Lambda_{n}|}\right]$
$\leq\beta\limsup_{n\to\infty}\frac{P\left[H_{\Lambda_{n},\emptyset}(\bullet)-H_{\Lambda_{n}}(\bullet)\right]}{|\Lambda_{n}|}+\left[\limsup_{n\to\infty}\frac{\log(Z_{\Lambda_{n},z,\beta})}{|\Lambda_{n}|}-P\left[\liminf_{n\to\infty}\frac{\log(Z_{\Lambda_{n},z,\beta,\bullet})}{|\Lambda_{n}|}\right]\right].$
@@
$\limsup_{n\to\infty}\frac{I(P_{\Lambda_{n}}\,|\,G_{\Lambda_{n},z,\beta})}{|\Lambda_{n}|}\leq 0$
which of course implies, by non-negativity of the relative entropy,
@@ 9.3. Minimizers are Gibbs measures
$\left\|G_{L,z,\beta,\zeta_{\Lambda_{k}}}-G_{L,z,\beta,\zeta}\right\|_{\text{TV}}$ $\leq 2\frac{\pi^{z}_{L}\left(|\mathrm{e}^{-\beta H_{L,\zeta}}-\mathrm{e}^{-\beta H_{L,\zeta_{\Lambda_{k}}}}|\right)}{\max\big\{Z_{L,z,\beta,\zeta},Z_{L,z,\beta,\zeta_{\Lambda_{k}}}\big\}}\leq 2\frac{\pi^{z}_{L}\left(2\cdot\mathds{1}_{B_{K}^{c}}\right)}{e^{-z\left\lvert L\right\rvert}}$
$=4\mathrm{e}^{z\left\lvert L\right\rvert}\,\pi^{z}_{L}\left(B_{K}^{c}\right),$
@@
Now note that, if $0=\delta F(P)=\lim_{n\to\infty}\frac{1}{\left\lvert\Lambda_{n}\right\rvert}I(P_{\Lambda_{n}}\,|\,G_{\Lambda_{n},z,\beta})$, then, by the translation invariance of $P$, it follows that $P_{\Lambda}\ll G_{\Lambda,z,\beta}$ with some density $g_{\Lambda}$ for large enough cubes (not necessarily centered at the origin) $\Lambda$. For every subset $\Delta\subset\Lambda$, we consider the restriction
@@
Furthermore, using $\Delta(l)=\Delta(l+1)\setminus\Lambda_{k}(l+1)$ for $0\leq l\leq m^{d}-1$,
@@
[…]
@@
$\int\,P(\mathrm{d}\zeta)\,f(\zeta)$ $\stackrel{{\scriptstyle(1)}}{{=}}\int\,G_{\Lambda,z,\beta}(\mathrm{d}\zeta)\,g_{\Lambda,\Delta}(\zeta)\,f(\zeta)$
@@
$\stackrel{{\scriptstyle(3)}}{{=}}\int\,G_{\Lambda,z,\beta}(\mathrm{d}\zeta)\,g_{\Lambda,\Delta\setminus L}(\zeta)\,\left[\int\,G_{L,z,\beta,\zeta}(\mathrm{d}\omega)\,f(\omega_{L}\zeta_{L^{c}})\right]$
@@
$\stackrel{{\scriptstyle(5)}}{{=}}\int\,P(\mathrm{d}\zeta)\,\left[\int\,G_{L,z,\beta,\zeta_{\Lambda_{k}}}(\mathrm{d}\omega)\,f(\omega_{L}\zeta_{\Lambda_{k}\setminus L})\right]$